# The nature of ASASSN-24fw’s occultation: modelling the event as dimming by optically thick rings around a sub-stellar companion

### v1

@@ 4.1.2 Occulting disc model
We now understand that the dimming is not cause by a single solid body but by a system with large area, varying density and optical thickness. So in the next step, we try to model the lightcurve by a point like physical body having a massive circum-planetary disc similar to that of Aur type systems (for example, Rodriguez et al. (2015, 2016); Rappaport et al. (2019b); van der Kamp et al. (2022). We call this model as the Occulting disc model. To compute this model, we use the Python package named pyPplusS333https://github.com/EdanRein/pyPplusS (Rein & Ofir, 2023). In our model, the dimming is caused by elliptical projection of the massive circum-planetary rings on to the star. Unlike the short dips caused by transiting planets which cover a small portion of the stellar surface, the large scale discs block a significant portion of the stellar area and hence require a different approach to calculate the stellar brightness during dimming. pyPplusS’s approach is to use the Polygon-Segments algorithm which efficient and stable algorithm.
@@
Typically, the parameters for computing a model in pyPplusS are: planet radius (Rp), the inner diameter of the ring (Rin), the outer diameter of the ring (Rout), the inclination angle of the ring from the orbital plane of the exoplanet (), the azimuthal angle by which the ring is rotated about the z-axis (), opacity (), and the limb-darkening coefficients of the star. In addition, the code requires time dependent coordinates of the exoplanet’s center and . We modified our modelling approach by considering a point like exoplanet surrounded by a massive ring system with sharp edges which is comparable to or larger than the stellar size. We further simplify our model by considering = (), and = dimensionless impact parameter ( 0). Here is the transverse velocity of the system which is assumed constant throughout the event, and is the time of the closest approach. Our simplification is made in a way that the exoplanet is traversing the star in an edge-on orbit and across its diameter. Thus our model parameters are: , , , , , and .
@@
We explored this parameter space by looking for models that could produce the 275 day long transit event, while also minimizing the values. Here also we have made use of the Python package . Obviously, the models that would produce such long transits need to have low transverse velocities that likely indicate the large semi-major axis between the occulter and the primary star. Our best-fitted model is shown in the lower panel of Figure (4) and the parameters with their 16th, 50th, and 84th percentile values are shown in Table (5). We see that our model fits the lightcurve well, but does not accurately describe the lightcurve because of a couple of things. First, the assumption that we made that rings have sharp edges is not entirely correct, and second, the trajectory may not be traversing through the diameter of the star. Nevertheless, our model gives an estimate on the geometry of the system. We see that our planet is surrounded by massive rings with the outer diameter being roughly 16R or 0.17 au. This system is inclined to the orbital plane by nearly 40o, rotated about the rotational axis by nearly 50o and the transverse velocity of the occulter is about 1R per day or nearly 17 km/s.
@@
As a check, we try another method to compute the transverse velocity of the occulter available in the literature. For example, van Werkhoven et al. (2014) calculate the transverse velocity based on the maximum gradient of the lightcurve. It requires the stellar linear limb-darkening parameter (), of the star. This parameter was computed using the jktld code of Southworth (2015). This code uses the stellar model parameters of the star to linearly interpolate the tables from Sing (2010) and calculates u for the required band passes. For ASASSN-24fw, we obtain u = 0.48 for the g’-band data of ASAS-SN survey. Using this value of , the steepest time gradient of the light curve, , and taking R= (from Table (2)), we get an estimate on :
@@
For ASASSN-24fw, we get the maximum value of the gradient of the lightcurve is 4.970.43 10-6 Lday-1. Therefore, is 13.74 2.49 km/s. This is close to the orbital speed of 11.5 km/s derived from Kepler’s third law and assuming an orbital period of 16000 days (Nair & Denisenko, 2024) and a orbital semi-major axis of 17 au. This is also not very different from the velocity that we get from our lightcurve model. Now, we can estimate the radius of the occulter () = x /2 0.16 0.02 au. This is similar to the outer radius of the disc that we get from our model 0.17 au. Since for a stable disc, its radius should be smaller than the Hill radius of the secondary, the lower limit we get on the mass of the secondary is 3()3> 3.42 . Therefore, we can safely assume that the secondary is most compatible with a super-Jupiter or a brown dwarf with ultra-massive rings.
@@ 5.1 Is ASASSN-24fw an R Corona Borealis (R CrB) type variable?
R CrB stars also show semi-regular pulsations with periods of multiple tens of days and amplitudes up to a few tenths of a magnitude (Percy, 2023); in contrast, as mentioned before, the star does not show significant variability in its TESS light curves. R CrB type variables come under the parent class of hydrogen-deficient stars, are found in a temperature range similar to that of ASASSN-24fw, and are known to show similar long (although sporadic) dimming behaviour (Clayton, 1996, 2012; Schaefer, 2024; Crawford et al., 2025). R CrB stars are also known to exhibit mid-IR excess due to warm circum-stellar dust shells (Chesneau et al., 2014; Rao & Lambert, 2015; Tisserand et al., 2020). The spectrum of R CrB stars are known for their hydrogen-less erratic and asymmetric dimming profiles and are dominated by mostly helium, nitrogen, and carbon enrichment in their atmospheres (see (Asplund et al., 2000; Pandey et al., 2021)). We found deep hydrogen absorption lines and no helium emission or P Cygni profile in our Keck optical spectrum (the second plot in Figure (2)). Moreover, no carbon absorption features or other spectroscopic R CrB signatures were found in our optical spectrum (see Pandey et al. 2021). As the available spectra were taken during the eclipse, we have not performed a detailed spectral analysis in search for other elements characteristic of hydrogen deficient stars (eg. Germanium Saini & Pandey 2025). However, this will be performed later when once we obtain a post-egress spectrum. Nevertheless it is highly unlikely that ASASN-24fw is a R CrB type of variable.
@@ 5.2 Is ASASSN-24fw a YSO?
YSOs are usually found near molecular clouds and often show optical variability and accretion signatures due to high magnetic fields and active circum-stellar zone comprised of an protoplanetary/proto-stellar material around them (Joy, 1945; Gahm et al., 2008; Zsidi et al., 2025). The circum-stellar material around a T Tauri type of variable star is divided in to two components: (1) the inner component, which is warmer and contains finer dust that accretes on to the host and (2) the outer component which is mixture of gaseous and solid material. Together they may manifest as two solid bodies causing infrared excess in the SED (eg. Varga et al. 2018 and references within). In the case of ASASSN-24fw, the persistent infrared excess is also fit by two black bodies.
@@
Although the shape of the SED and the mass of ASASSN-24fw is similar to that of a T Tauri star, in the only pre-dimming spectrum that we have from Gaia DR3, no H emission line appears which may point towards any accretion activity. Similarly, in our low resolution spectrum obtained from Keck, we do not see any emission signatures. In our Keck high-resolution spectrum taken in January 2025, we see several deep hydrogen absorption lines. If the Ca II HK lines are unaffected by the occulter, then their strength resembles to that of an F-type main-sequence star. However, in the same spectrum, we see that the H absorption line is shallower unlike other Hydrogen series lines and has a shape. If the shape is not due to spectral noise then a component of the occulting material is Hydrogen rich and re-emitting the incident stellar radiation into H region. Moreover, the otherwise flat lightcurve of ASASSN-24fw only shows deep occultation every 43.8 years (Nair & Denisenko, 2024) that is likely caused by a large solid object. This is unlike any other lightcurve of a T Tauri star where we expect regular episodes of irregular dips.
@@
In most cases the circum-stellar discs around YSOs are bright in X-ray and also radio sources owing to the strong magnetic fields and high temperatures from ongoing accretion activity (Feigelson & Montmerle, 1999; Espaillat et al., 2019). We searched through known X-ray catalogs like ROSAT 2RXS (Boller et al., 2016), XMMDR10 (Webb et al., 2022), XMMSL2 (XMM-SSC, 2018), Chandra CSC2 (Evans et al., 2024), and swift2SXPS (Kosiba et al., 2024) and found no pointing at the location of ASASSN-24fw. Thus we do not have evidence that the companion of ASASSN-24fw is a compact binary emitting X-rays. Given the lack of erratic photometric variability, no X-ray detections, and no evidence of narrow emission lines in the optical spectrum at the time of dimming, we do not believe that the dimming of ASASSN-24fw results from the activity in an accretion disc or that ASASSN-24fw is surrounded by a compact star. As mentioned earlier, the peculiar velocity of ASASSN-24fw indicates that it does not appear to have interacted with any star forming regions at least in past 50 Ma. Further, the peculiar velocity indicates that it belongs to the older population of stars in thick disc. Therefore the origin of the infrared excess is unlikely due to proto-planetary material.
5.3 Is the occultation caused by circum-stellar material?
As mentioned earlier, many young stars show the "dipping" behavior (Cody et al., 2014; Ansdell et al., 2016, 2019), i.e., multiple, repeated, short, and irregular dips in brightness. The most common known dipper stars are also young main sequence stars of F- or G- spectral type, but still accreting like classical T Tauri stars. For these accreting stars, Ansdell et al. (2016), and Nuñez et al. (2017) show a positive correlation between the depth of dips and the amount of infrared excess. In ASASSN-24fw we see an infrared excess as well as a deep but long dimming causing a flat base unlike the irregular shaped lightcurve by "dippers". The circum-stellar material causing infrared excess appears to have spread out rather than concentrated in clumps. This is supported by the fact that the mid-infrared NEOWISE lightcurve is also non-variable. Now the question arises whether this material accretes on to the star and is responsible for the ’w’ shaped H-alpha line that we see.
From the lightcurve we also see a phenomenon called as "forward-scattering", meaning the post-dimming color is not very different or slightly bluer than the original color of the star. This is possible when the material contains bodies that block the starlight but do not cause extinction effects. In terms of our Solar System, strong forward scattering has been observed in the Saturn B-ring system which is discussed by Hedman & Stark (2015). So the fact that we see forward scattering in the post-dimming lightcurve of the ASASSN-24fw could hint towards a dense mixture of dust and gas. This is also what we inferred from our dust composition analysis based on LCOGT data using optool. However, large icy bodies also cannot be ruled out. This composition with the large size of the disc suggests that the system is young ( 1 Ga), as the dust is usually depleted over time by radiation pressure, Poynting-Robertson drag, or accretion.
@@ 5.4 Nature and age of the system
The longer orbital period, depth of dimming and eclipse duration of ASASSN-24fw makes it a unique, and observationally challenging binary system. The crude stellar analysis based on our available data suggests that the star is unlikely to be a YSO although it has a persistent infrared excess. It is likely that the star has an active planetary environment and the catastrophic collisions between planetesimals has caused a debris disc around it. Although this can be verified, the distance of 1kpc would make it challenging to resolve the discs. Nevertheless, mid-infrared studies from JWST can tell us about the composition of the circum-stellar environment. Follow up series of observations from NIRCam or VLT SPHERE can infact be useful to measure the proper motion and companionship of the companion that we found in our LCOGT images.
@@
One of the distinguishing factors between the super-Jupiters and brown dwarfs can be their discs. If the secondary is a super-Jupiter (with 5–13 MJup), it could host a large disc, provided it is young or resides in a debris-rich environment. Such a system could persist, especially at 20 au where dynamical clearing is slow (for example, Liu et al. (2023)). However, forming a super-Jupiter hosting circum-planetary disc with a size of 272 times the size of Jupiter is complicated as it can become gravitationally unstable owing to reduced lifetime. Instead, the companion being a more massive brown dwarf is more likely scenario.
Brown dwarfs are believed to form like stars via a gravitational collapse, and thus can retain the surrounding material (Stamatellos & Whitworth, 2009; Stamatellos & Herczeg, 2015), especially when the disc radius is well outside the Roche limit of the star. A disc of 0.20 au radius is consistent with observed discs around young brown dwarfs (e.g., in star-forming regions like Taurus Kurtovic et al. 2021). Such discs comprise of gas and fine dust (micron-sized grains) that scatter or absorb the stellar light effectively. This material is possibly replenished, from the ongoing accretion (Stamatellos & Herczeg, 2015; Zhu, 2015).
Although replenishing process may be responsible for the observed disc, this process is efficient only in the first few million years of the formation. As the system evolves, various "outside" forces come affect the process. Thus, if we consider that the star is an older main-sequence star, then an age of about 2 Ga introduces new challenges for the disc survival. On such longer timescales, the brown dwarf’s disc is expected to be depleted due to dynamical and radiative processes. Moreover, as a mature disc looses its opacity due to dust coagulation or dispersal, a significant luminosity drop in ASASN-24fw due to a older disc seems improbable. Although at larger separations, the brown dwarf’s disc would experience minimal stellar radiation, but the internal processes (e.g., viscous evolution) and external perturbations (e.g., passing stars) would likely deplete it long before 2 Ga. Therefore, a massive, opaque circum-planetary disc that is surviving for so long on its own is unlikely unless it is replenished by some process (e.g. by accretion from the circum-stellar environment which could be rich in material due to active collisions). Another process for the replenishment is the Bondi–Hoyle–Lyttleton accretion (Hoyle & Lyttleton, 1939; Bondi & Hoyle, 1944) which would have played a role if the brown dwarf would have passed through an unusually dense region of the ISM (e.g., a molecular cloud or diffuse gas) in the past.

### v2

@@ 4.1.2 Occulting disc model
We now understand that the dimming is not cause by a single solid body but by a system with large area, varying density and optical depth. So in the next step, we try to model the lightcurve by a point like physical body having a massive circumplanetary disc similar to that of Aur type systems (for example, (Rodriguez et al., 2015, 2016; Rappaport et al., 2019a; van der Kamp et al., 2022). We call this model the Occulting disc model. To compute this model, we use the Python package named pyPplusS444https://github.com/EdanRein/pyPplusS (Rein and Ofir, 2023). In our model, the dimming is caused by elliptical projection of the massive circumplanetary rings on to the star. Unlike the short dips caused by transiting planets which cover a small portion of the stellar surface, the large scale discs block a significant portion of the stellar area and hence require a different approach to calculate the stellar brightness during dimming. pyPplusS’s approach is to use the Polygon-Segments algorithm which efficient and stable algorithm.
@@
Typically, the parameters for computing a model in pyPplusS are: planet radius (), the inner diameter of the ring (), the outer diameter of the ring (), the inclination angle of the ring from the orbital plane of the exoplanet (), the azimuth angle by which the ring is rotated about the z-axis (or the exoplanet’s spinning axis) (), opacity (), and the limb-darkening coefficients of the star. We also consider a uniform source and do not find any significant difference in the generated model. In addition, the code requires time dependent coordinates of the exoplanet’s center and . We modified our modelling approach by considering a point like exoplanet surrounded by a massive ring system with sharp edges which is comparable to or larger than the stellar size. We further simplify our model by considering = (), and = impact parameter (or b 0). Here is the transverse velocity of the system which is assumed constant throughout the event, and is the time of the closest approach. Our simplification is made in a way that the exoplanet is traversing the star in an edge-on orbit and across its diameter. Thus our model parameters are: , , , , , and .
@@
We explored this parameter space by looking for models that could produce the 275 day long transit event, while also minimizing the values. Here also we have made use of the Python package emcee. Obviously, the models that would produce such long transits need to have low transverse velocities that likely indicate the large semi-major axis between the occulter and the primary star. Our best-fitted model is shown in the lower panel of Figure (4) (uniform source in green color and limb-darkened source in red color) and the fitted parameters with their 16th, 50th, and 84th percentile values are shown in Table (5). We see that our model fits the lightcurve well, but does not accurately describe the lightcurve because of a couple of things. First, the assumption that we made that rings have sharp edges is not entirely correct, and second, the trajectory may not be traversing through the diameter of the star. Nevertheless, our model gives an estimate on the geometry of the system. We see that our planet is surrounded by optically thick rings with the outer diameter being roughly 16 or 0.17 au. This system is inclined to the orbital plane by nearly , rotated about the rotational axis by nearly and v(t) is about 1 per day or 17 km/s. We note that in practice v(t) is expected be degenerate with other model parameters. However, in our simplistic approach, we have constrained our model by fixing the size of the exoplanet to a point and assuming an edge-on trajectory traversing the diameter. This constrains the parameter space to be explored and resolves the degeneracies. Moreover, the orbit of the planet may be inclined at an angle. But the projection of the rings on the star blocks an area of the star equal to a portion of an ellipse. This area is proportional to the amount and duration of the light getting blocked. Since this is a massive and long duration dimming, the area of the star blocked is much larger than the ones blocked by a single transiting exoplanet. So we are not modelling the transit by an exoplanet but the transit by projected rings that are blocking the stellar light.
@@
As a check on our computed v(t) using lightcurve model, we try another method to compute the transverse velocity of the occulter. For example, van Werkhoven et al. (2014) calculate the transverse velocity based on the maximum gradient of the lightcurve. It requires the stellar linear limb-darkening parameter (), of the star which can be computed using the jktld code of Southworth (2015). This code uses the stellar model parameters of the star to linearly interpolate the tables from Sing (2010) and calculates u for the required band passes. For ASASSN-24fw, we obtain u = 0.48 for the g’-band data of ASAS-SN survey. Using this value of , the steepest time gradient of the light curve, , and taking R= (from Table (2)), we can get another estimate on :
@@
From the lightcurve, we get the maximum value of the gradient = 4.970.43 10-6 Lday-1, (where L is normalized to L). Therefore, we obtain 13.74 2.49 km/s. This is close to the orbital speed of 11.5 km/s derived from Kepler’s third law, considering an orbital period of 16000 days (Nair and Denisenko, 2024) and an orbital semi-major axis of 17 au. This is also not very different from the velocity that we get from our lightcurve model. Then the occulter radius () = /2 0.16 0.02 au. This is similar to the outer radius of the disc that we get from our model 0.17 au. Since for a stable disc, its radius should be smaller than the Hill radius of the secondary, the lower limit we get on the mass of the secondary is 3()3> 3.42 . Therefore, we can safely assume that the secondary is most compatible with a super-Jupiter or a brown dwarf with optically thick rings.
4.2 Extinction calculations using LCOGT data
Our aim of obtaining multi-band LCOGT observations was to monitor the variation in color of ASASSN-24fw during dimming. The variation in color can then be used to compute the specific reddening vector and hence the composition of the occulting material. Unfortunately, our post-dimming baseline of ASASSN-24fw using LCOGT observations in the four SDSS filters is not very long. However, it is still enough to compute the reddening vector and hence constrain the primary composition of the occulting material. We measure mean colors for ASASSN-24fw of - = 0.40 0.07 mag, - = 0.16 0.05 mag, and - = 0.28 0.10 mag. For comparison, we estimate the ISM extinction using the Python package extinction (Barbary, 2021) to extinguish the stellar photosphere model using the Fitzpatrick ’99 (Fitzpatrick, 1999) extinction law and obtain synthetic photometry in the necessary filter bands using the Python package pyphot (Fouesneau, 2025). We find that the photospheric colors for ASASSN-24fw should thus be - = 0.42 mag, - = 0.25 mag, and - = 0.17 mag, assuming a distance of 1044 pc. The observed reddening is therefore not consistent with ISM extinction, being significantly stronger at longer wavelengths than expected.
We generate color-color plots using , , and bands and use optool (Dominik et al., 2021) to calculate the opacities (= +) values for a range of dust grain compositions and two size distributions. We used the built-in refractive indices from optool to generate the values for materials including astronomical silicate (Draine, 2003), various crystalline and amorphous iron-magnesium silicates (Dorschner et al., 1995; Jaeger et al., 1998; Fabian et al., 2001; Suto et al., 2006), carbon-bearing species (Draine and Malhotra, 1993; Henning and Stognienko, 1996), quartz (Kitamura et al., 2007), and iron particles (Henning and Stognienko, 1996).
We consider a log-normal grain distribution with a width of 0.5 dex and a free peak size, considering contributions from sizes between 0.1 and 1.1m in steps of 0.01m when calculating , determined from the complex refractive indices assuming Mie theory (Mie, 1908). For a Sun-like star, the typical blowout size for a dust grain is 0.5m (Krivov, 2010), so we include transient dust with this range of grain sizes. The optical depth and extinction of a dust cloud that fully covers the stellar disc are then calculated from the values for each combination of composition and size distribution, which are then compared to the measured reddening vectors by least-squares fit to identify those which best match the observations. For ASASSN-24fw, we find that no combination of compositions available in optool and grain size produces a reddening vector that fits all the data. We can match individual sizes and compositions to separately fit the bulk of measurements in the - vs - and - vs - color-color plots but no single material adequately reproduces the reddening vectors in both. Forés-Toribio et al. (2025) found silicate materials around 1 m in size could reproduce the - opacity gradient; we find astronomical silicate dust grains around 0.4 to 0.6 m in size is the best, albeit imperfect, overall match to the reddening.
@@ 5.1 Is ASASSN-24fw an R Corona Borealis (R CrB) type variable?
R CrB stars also show semi-regular pulsations with periods of multiple tens of days and amplitudes up to a few tenths of a magnitude (Percy, 2023); in contrast, as mentioned before, the star does not show significant variability in its TESS light curves. R CrB type variables come under the parent class of hydrogen-deficient stars, are found in a temperature range similar to that of ASASSN-24fw, and are known to show similar long (although sporadic) dimming behaviour (Clayton, 1996, 2012; Schaefer, 2024; Crawford et al., 2025). R CrB stars are also known to exhibit mid-IR excess due to warm circumstellar dust shells (Chesneau et al., 2014; Rao and Lambert, 2015; Tisserand et al., 2020). The spectrum of R CrB stars are known for their hydrogen-less erratic and asymmetric dimming profiles and are dominated by mostly helium, nitrogen, and carbon enrichment in their atmospheres (see Asplund et al., 2000; Pandey et al., 2021). We found deep hydrogen absorption lines and no helium emission or P Cygni profile in our Keck HiRES optical spectrum (the second plot in Figure (2)) and also the post-dimming Magellan MIKE Spectrum (See Figure (5)). Moreover, no carbon absorption features or other spectroscopic R CrB signatures are seen in our optical spectrum (see Pandey et al., 2021). In this study, we have not performed a detailed spectral analysis in search for other elements characteristic of hydrogen deficient stars (eg. Germanium Saini and Pandey 2025). However, this will be performed later when once we obtain a post-egress spectrum. Nevertheless it is highly unlikely that ASASSN-24fw is a R CrB type of variable.
@@ 5.2 Is ASASSN-24fw a YSO?
YSOs are of F- or G- spectral type, usually found near molecular clouds, and often show optical variability and accretion signatures due to high magnetic fields and active circumstellar zone comprised of an protoplanetary/proto-stellar material around them (Joy, 1945; Gahm et al., 2008; Zsidi et al., 2025). All YSOs have infrared excess and the ones that show optical variability are called as T Tauri type of variables (or "dippers"). As discussed earlier, this variability is manifested as irregular and non-periodic dips and is linked with clumpy circumstellar material (Cody et al., 2014; Ansdell et al., 2016, 2019). It has been found by Ansdell et al. (2016), and Nuñez et al. (2017), that a positive correlation exists between the depth of dips and the amount of infrared excess.
The circumstellar material around a T Tauri type of variable star is divided in to two components: (1) the inner component, which is warmer and contains finer dust that accretes on to the host and (2) the outer component which is mixture of gaseous and solid material. Together they may manifest as two solid bodies causing infrared excess in the SED similar to what we see here (e.g. Varga et al. 2018 and references within). In ASASSN-24fw we see an infrared excess as well as a deep but long dimming causing a flat base unlike the irregular shaped lightcurve. Therefore the circumstellar material causing infrared excess in ASASSN-24fw appears to have spread out rather than concentrated in clumps. This is supported by the fact that the mid-infrared NEOWISE lightcurve is also non-variable.
@@
Although the shape of the SED and the mass of ASASSN-24fw is similar to that of a T Tauri star, in the only pre-dimming low-resolution spectrum that we have from Gaia DR3, no H emission line appears which is an indicator of accretion activity. Similarly, in the low resolution during-dimming spectrum obtained from Keck, we do not see any emission signatures. In our Keck high-resolution spectrum taken in January 2025, we see several deep hydrogen absorption lines. If the Ca II HK lines are unaffected by the occulter, then their strength resembles to that of an F-type main-sequence star. However, in the same spectrum, we see that the H absorption line is shallower unlike other hydrogen series lines and has a shape resembling a central emission feature (as already discussed).
If the shape is not due to spectral noise then a component of the occulting material is Hydrogen rich and re-emitting the incident stellar radiation into H region. The otherwise flat lightcurve of ASASSN-24fw only shows deep occultation every 43.8 years (Nair and Denisenko, 2024) which indicates that the occulter is large and optically thick. This is unlike any other lightcurve of a T Tauri star where we expect regular episodes of irregular dips.
@@
In most cases the accreting circumstellar discs around YSOs are bright in X-ray and radio wavelengths owing to the strong magnetic fields and high temperatures from ongoing accretion activity (Feigelson and Montmerle, 1999; Espaillat et al., 2019). If the companion is a stellar remnant, we can also see X-ray emission from the system. We searched through known X-ray catalogs like ROSAT 2RXS (Boller et al., 2016), XMMDR10 (Webb et al., 2022), XMMSL2 (XMM-SSC, 2018), Chandra CSC2 (Evans et al., 2024), and swift2SXPS (Kosiba et al., 2024) and found no pointing at the location of ASASSN-24fw. Thus we do not have evidence that the ASASSN-24fw has a compact companion or there are X-rays emitted from accretion acitivity.
In Appendix (A), we show a high-resolution post-dimming spectrum taken on 13 October 2025. Here we can see a deep H line in absorption and not in emission. If ASASSN-24fw were an YSO, and the mid-infrared excess causing material was a proto-planetary material, we would expect strong H emission line. It may be possible that accretion may be happening in a different location on the stellar surface and not along our line of sight. However, we do not see any reason as to why this accretion spot has to change post-dimming (as we have already rules out fast stellar rotation). This, and the absence of Li line, the age indicator for main sequence stars (Llorente de Andrés et al., 2021), supports the fact that the infrared excess causing material is not proto-planetary and is likely to be a debris disc. Also as mentioned earlier, the peculiar velocity of ASASSN-24fw indicates that it does not appear to have interacted with any star forming regions at least in past 50 Ma and that it belongs to the older population of stars in thick disc. These things support the fact that ASASSN-24fw is unlikely to be an YSO.
@@ 5.3 Nature and age of the system
The larger distance, longer orbital period, large depth of dimming and longer eclipse duration of ASASSN-24fw makes it a unique, and observationally challenging system. The crude stellar analysis based on our available data suggests that the star is unlikely to be a YSO although it has a persistent infrared excess. It is likely that the star has an active planetary environment and the catastrophic collisions between planetesimals has caused a debris disc around it. Although this can be verified, the distance of 1kpc would make it challenging to resolve the discs. Nevertheless, mid-infrared studies from JWST can tell us about the composition of the circumstellar environment. Follow up series of observations from NIRCam or VLT SPHERE can infact be useful to measure the proper motion and confirm the companionship of the faint source that we found in our LCOGT images.
@@
One of the distinguishing factors between the super-Jupiters and brown dwarfs can be their discs. If the secondary is a super-Jupiter (with 5–13 MJup), it could host a large disc, provided it is young or resides in a debris-rich environment. Such a system could persist, especially at 20 au where dynamical clearing is slow (for example, Liu et al. (2023)). However, forming a super-Jupiter hosting a optically thick disc with a size of 272 times the size of Jupiter is complicated as it can become gravitationally unstable owing to reduced lifetime. Instead, the companion being a more massive brown dwarf is more likely scenario.
Brown dwarfs are believed to form like stars via a gravitational collapse, and thus can retain the surrounding material (Stamatellos and Whitworth, 2009; Stamatellos and Herczeg, 2015), especially when the disc radius is well outside the Roche limit of the star. A disc of 0.20 au radius is consistent with observed discs around young brown dwarfs (e.g., in star-forming regions like Taurus (Kurtovic et al., 2021)). Such discs comprise of gas and fine dust (micron-sized grains) that scatter or absorb the stellar light effectively. This material is possibly replenished, from the ongoing accretion (Stamatellos and Herczeg, 2015; Zhu, 2015).
Although replenishing process may be responsible for the observed disc, this process is efficient only in the first few million years of the formation. As the system evolves, various "outside" forces come affect the process. Thus, if we consider that the star is an older main-sequence star, then an age of about 2 Ga introduces new challenges for the disc survival. On such longer timescales, the brown dwarf’s disc is expected to be depleted due to dynamical and radiative processes. Moreover, as a mature disc looses its opacity due to dust coagulation or dispersal, a significant luminosity drop in ASASSN-24fw due to a older disc seems improbable. Although at larger separations, the brown dwarf’s disc would experience minimal stellar radiation, but the internal processes (e.g., viscous evolution) and external perturbations (e.g., passing stars) would likely deplete it long before 2 Ga. Therefore, a massive, opaque circumplanetary disc that is surviving for so long on its own is unlikely unless it is replenished by some process (e.g. by accretion from the circumstellar environment which could be rich in material due to active collisions).